Canadian Situation

The programme has identified three priorities

• SNO - most of it

• ATLAS - half of HCAL and FCAL + electronics

• ISAC - radioactive beam facility at TRIUMF

Long range plan two years ago concluded

• Couldn’t afford BaBar or CDF

• NSERC Grant Selection Committee approved mini-BaBar • Recognized the need to fund young groups

CDF group has been proceeding with Run II proposal

• Have several new people (ATLAS) involved – Bob Orr (Canadian ATLAS spokesperson)

– John Mayer (Senior ATLAS RA)

Run II Plans

Now have group of 14 faculty, RA’s and students

• Focus on SVX II mounts, inchworms & alignment

• Also getting involved in offline software (again)

Have proposed equipment and 3 years of operating support

• Would bring in $250K of equipment

• Similar level of operating support

• Also some University funding Will know outcome in March 1998

• Cautiously optimistic

• Not expecting significant growth in CDF effort